# Provable In-Context Learning of Nonlinear Regression with Transformers

### v1

@@ C.1 Proof of Proposition˜2
Define $u_{k}=K(p_{k}-\delta)$ and $U_{k}=K(p_{k}+\delta)$. Then, from the concentration property (see (9)), $|\mathcal{V}_{k}|\in[\frac{u_{k}}{K}N,\frac{U_{k}}{K}N]$ for constants $u_{k},U_{k}=\Theta(1)$. Therefore,
@@
By the concentration condition in (9), $|\mathcal{V}_{k}|\geq\frac{u_{k}}{K}N$ for some $u_{k}=\Theta(1)$, and $N/|\mathcal{V}_{k}|=\Theta(1/\delta)$ (since $\delta=o(1)$ is the imbalance parameter). Thus,
@@
where the last equality follows because $1/u_{k}-1/K=\Theta(\delta)$ from (9).
@@ C.2 Proof of Proposition˜3
where the last equality uses $\frac{1}{u_{k}}-\frac{1}{K}=\Theta(\delta)$ from (9). Thus, $1-\mathrm{Attn}_{k}^{(t)}=\mathcal{O}(\delta)$.
@@
C.3 Proof of Theorem˜1
Recall from Lemma˜2 and its proof that the prediction error $\mathcal{L}(\theta)$ defined in (7) can be expressed as
$\mathcal{L}(\theta)=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(t)})^{2}\Theta(L^{2}\Delta^{2})\right],$
@@ C.3 Proof of Theorem˜1
$\mathcal{L}(\theta)$ $=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(T_{f}^{*}+1)})^{2}\Theta(L^{2}\Delta^{2})\right]$
@@
Appendix D Proofs for Sharp $L$-Regime
D.1 Proof of Proposition˜4
@@
D.2 Proof of Theorem˜2
As in the proof of Theorem˜1, the prediction error $\mathcal{L}(\theta)$ (from (7)) can be written as
$\mathcal{L}(\theta)=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(t)})^{2}\Theta(L^{2}\Delta^{2})\right].$
@@ D.2 Proof of Theorem˜2
$\mathcal{L}^{(T_{s}^{*})}(\theta)$ $=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(T_{s}^{*})})^{2}\Theta(L^{2}\Delta^{2})\right]$
@@
Under the scaling regime for sharp $L$, either $\delta=o(1)$ or $\epsilon=o(1)$ (since the in-context learning regime assumes both go to zero), and hence $\mathcal{L}^{(T_{s}^{*})}(\theta)=\mathcal{O}(\epsilon^{2})$ as required.
This completes the proof of Theorem˜1.
Appendix E Proof of Proposition˜1
@@
where the last inequality is based on the non-degenerate $L$-Lipschitz in (4).

### v2

@@ D.1 Proof of Proposition˜2
Define $u_{k}=K(p_{k}-\delta)$ and $U_{k}=K(p_{k}+\delta)$. Then, from the concentration property (see Eq.˜9), $|\mathcal{V}_{k}|\in[\frac{u_{k}}{K}N,\frac{U_{k}}{K}N]$ for constants $u_{k},U_{k}=\Theta(1)$. Therefore,
@@
By the concentration condition in Eq.˜9, $|\mathcal{V}_{k}|\geq\frac{u_{k}}{K}N$ for some $u_{k}=\Theta(1)$, and $N/|\mathcal{V}_{k}|=\Theta(1/\delta)$ (since $\delta=o(1)$ is the imbalance parameter). Thus,
@@
where the last equality follows because $1/u_{k}-1/K=\Theta(\delta)$ from Eq.˜9.
@@ D.2 Proof of Proposition˜3
where the last equality uses $\frac{1}{u_{k}}-\frac{1}{K}=\Theta(\delta)$ from Eq.˜9. Thus, $1-\mathrm{Attn}_{k}^{(t)}=\mathcal{O}(\delta)$.
@@
D.3 Proof of Theorem˜1
Recall from Lemma˜2 and its proof that the prediction error $\mathcal{L}(P;Q)$ defined in Eq.˜7 can be expressed as
$\mathcal{L}(P;Q)=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(t)})^{2}\Theta(L^{2}\Delta^{2})\right],$
@@ D.3 Proof of Theorem˜1
$\mathcal{L}(P;Q)$ $=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(T_{f}^{*}+1)})^{2}\Theta(L^{2}\Delta^{2})\right]$
@@
Appendix E Proofs for Sharp $L$-Regime
E.1 Proof of Proposition˜4
@@
E.2 Proof of Theorem˜2
As in the proof of Theorem˜1, the prediction error $\mathcal{L}(P;Q)$ (from Eq.˜7) can be written as
$\mathcal{L}(P;Q)=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(t)})^{2}\Theta(L^{2}\Delta^{2})\right].$
@@ E.2 Proof of Theorem˜2
$\mathcal{L}^{(T_{s}^{*})}(P;Q)$ $=\frac{1}{2}\sum_{k=1}^{K}\mathbb{E}\left[\mathds{1}\{x_{\mathrm{query}}=v_{k}\}(1-\mathrm{Attn}_{k}^{(T_{s}^{*})})^{2}\Theta(L^{2}\Delta^{2})\right]$
@@
Under the scaling regime for sharp $L$, either $\delta=o(1)$ or $\epsilon=o(1)$ (since the in-context learning regime assumes both go to zero), and hence $\mathcal{L}^{(T_{s}^{*})}(P;Q)=\mathcal{O}(\epsilon^{2})$ as required.
This completes the proof of Theorem˜2.
Appendix F Proof of Proposition˜1
@@
where the last inequality is based on the non-degenerate $L$-Lipschitz in Eq.˜4.